What is the Power of a rechargeable flashlight used by a lookout if it 

uses 60,000J for 50min?


1
Expert's answer
2021-01-12T12:10:38-0500

By the definition of the power, we have:


"P=\dfrac{W}{t}=\dfrac{60000\ J}{50\ min\cdot(\dfrac{60\ s}{1\ min})}=20\ W."

Answer:

"P=20\ W."